measure of interior angle of regular polygon is ten times measure of exterior angle. How many sides to polygon?
Answer by RAY100(1637)

Interior angle = (n-2)(180) / n
.
Exterior angle = 360/n
.
where n= number of sides
.
(n-2)(180)/n = 10 * 360/n
.
180n-360 = 3600
.
180n = 3960
.
n= 22
.
.
check ,,,interior = 20*180/22=163.636
.
exterior = 360 / 22 =16.3636,,,,,ok